# Analog In-memory Training on General Non-ideal Resistive Elements: The Impact of Response Functions

### v4

@@ Appendix G Proof of Theorem 1: Implicit Bias of Analog Training
Proof of Theorem 1 and 5.
We separately show that $\|\nabla f_{\Sigma}(W)\|\leq O((W^{\diamond}-W^{*})^{2})$ and $T(W^{S})\leq O((W^{\diamond}-W^{*})^{2})$.
Proof of $\|\nabla f_{\Sigma}(W)\|\leq O((W^{\diamond}-W^{*})^{2})$. The gradient of the penalized objective $f_{\Sigma}(W)$ is given by
$\nabla f_{\Sigma}(W)=\nabla f(W)+\Sigma\odot R(W).$ (49)
$\nabla f(W^{S})=$ $\ \nabla f(W^{*})+\nabla^{2}f(W^{*})(W^{S}-W^{*})+O((W^{S}-W^{*})^{2}),$ (50)
$\frac{G(W^{S})}{F(W^{S})}=$ $\ \frac{G(W^{\diamond})}{F(W^{\diamond})}+\nabla R(W^{\diamond})(W^{S}-W^{\diamond})+O((W^{S}-W^{\diamond})^{2}),$ (51)
we bound the gradient of the penalized objective as follows
$\ \|\nabla f_{\Sigma}(W)\|=\left\|\nabla f(W^{S})-\Sigma\odot\frac{G(W^{S})}{F(W^{S})}\right\|$ (52)
$=$ $\ \left\|\nabla^{2}f(W^{*})(W^{S}-W^{*})+O((W^{S}-W^{*})^{2})-\Sigma\odot(\nabla R(W^{\diamond})(W^{S}-W^{\diamond}))+O((W^{S}-W^{\diamond})^{2})\right\|$
$=$ $\ O((W^{S}-W^{*})^{2})+O((W^{S}-W^{\diamond})^{2})$
$=$ $\ O((W^{*}-W^{\diamond})^{2})$
where the last inequality holds by the definition of $W^{S}$.
Proof of $T(w^{S})\leq O((w^{\diamond}-w^{*})^{2})$. By the definition of effective update $T(W^{S})$, we have
$\ \left\|\frac{{\mathbb{E}}_{\xi}\left[U_{q}(W^{S},-\alpha\nabla f(W^{S};\xi))\right]-W^{S}}{\alpha}\right\|$ (53)
$=$ $\ \left\|{\mathbb{E}}_{\xi}[\nabla f(W^{S};\xi)]\odot F(W^{S})-{\mathbb{E}}_{\xi}[|\nabla f(W^{S};\xi)|]\odot G(W^{S})\right\|$
$\leq$ $\ \left\|\left(\nabla f(W^{S})-{\mathbb{E}}_{\xi}[|\nabla f(W^{S};\xi)|]\odot\frac{G(W^{S})}{F(W^{S})}\right)\right\|F_{\max}$
$\leq$ $\ \left\|\nabla f(W^{S})-{\mathbb{E}}_{\xi}[|\nabla f(W^{S};\xi)-\nabla f(W^{S})|]\odot\frac{G(W^{S})}{F(W^{S})}\right\|F_{\max}$
$\ +\left\|({\mathbb{E}}_{\xi}[|\nabla f(W^{S};\xi)|]-{\mathbb{E}}_{\xi}[|\nabla f(W^{S};\xi)-\nabla f(W^{S})|])\odot\frac{G(W^{S})}{F(W^{S})}\right\|F_{\max}$
$\leq$ $\ \|\nabla f_{\Sigma}(W)\|F_{\max}+\left\|({\mathbb{E}}_{\xi}[|\nabla f(W^{S};\xi)|]-{\mathbb{E}}_{\xi}[|\nabla f(W^{S};\xi)-\nabla f(W^{S})|])\odot\frac{G(W^{S})}{F(W^{S})}\right\|F_{\max}$
The first term in the RHS (RHS) of (53) is already bounded by (52). By inequality $||x|-|y||\leq|x-y|$ for any $x,y\in{\mathbb{R}}$, the second term in the RHS of (53) is bounded by
$\ \left\|({\mathbb{E}}_{\xi}[|\nabla f(W^{S};\xi)|]-{\mathbb{E}}_{\xi}[|\nabla f(W^{S};\xi)-\nabla f(W^{S})|])\odot\frac{G(W^{S})}{F(W^{S})}\right\|$ (54)
$\leq$ $\ \left\||\nabla f(W^{S})|\odot\frac{G(W^{S})}{F(W^{S})}\right\|$
$=$ $\ \left\||\nabla f(W^{S})-\nabla f(W^{*})|\odot\left(\frac{G(W^{S})}{F(W^{S})}-\frac{G(W^{\diamond})}{F(W^{\diamond})}\right)\right\|$
$\leq$ $\ O(|W^{S}-W^{*}|)\cdot O(|W^{S}-W^{\diamond}|)$
$=$ $\ O((W^{*}-W^{\diamond})^{2})$
Plugging back (52) and (54) into (53) shows $T(w^{S})\leq O((w^{\diamond}-w^{*})^{2})$. It is trivial to prove $T(W^{\diamond})=O(|W^{\diamond}-W^{*}|)$ and $T(W^{*})=O(|W^{\diamond}-W^{*}|)$ by the definition of $W^{S}$ and (50). ∎
@@ Appendix H Proof of Theorem 2: Convergence of Analog SGD
Define the saturation vector $R(W_{k})\in{\mathbb{R}}^{D}$ by
$R(W_{k}):=$ $\ {F(W_{k})^{\odot 2}-G(W_{k})^{\odot 2}}={(F(W_{k})+G(W_{k}))\odot(F(W_{k})-G(W_{k}))}$ (59)
@@
$=$ $\ -\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{R(W_{k})}\leq 0.$
@@
$=$ $\ -\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{R(W_{k})}+\frac{\alpha\sigma^{2}}{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$
@@
$\leq$ $\ f(W_{k})-\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{R(W_{k})}+\alpha^{2}LF_{\max}^{2}\sigma^{2}+\frac{\alpha\sigma^{2}}{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$
@@
$\mathbb{E}_{\xi_{k}}[f(W_{k+1})]\leq$ $\ f(W_{k})-\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{R(W_{k})}+\alpha^{2}LF_{\max}^{2}\sigma^{2}+\alpha\sigma^{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$ (66)
@@
$E^{\texttt{ASGD}}_{K}=$ $\ \frac{1}{K}\sum_{k=0}^{K}{\mathbb{E}}[\|\nabla f(W_{k})\odot F(W_{k})+|\nabla f(W_{k})|\odot G(W_{k})\|^{2}+4\|\nabla f(W_{k})\|^{2}_{R(W_{k})}]$ (67)
@@
Assuming the saturation never happens during the training, i.e. $R(W_{k})\geq R^{\texttt{RL}}_{\min}>0$ for all $k\in[K]$, we get a tighter bound in (66) by leveraging $\left\|\nabla f(W_{k})\right\|^{2}_{R(W_{k})}\geq\min\{R(W_{k})\}\left\|\nabla f(W_{k})\right\|^{2}\geq R^{\texttt{RL}}_{\min}\left\|\nabla f(W_{k})\right\|^{2}$
$\mathbb{E}_{\xi_{k}}[f(W_{k+1})]\leq$ $\ f(W_{k})-\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{R(W_{k})}+\alpha^{2}LF_{\max}^{2}\sigma^{2}+\alpha\sigma^{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$ (68)
@@
$=$ $\ \frac{4F_{\max}^{2}}{R^{\texttt{RL}}_{\min}}\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L}{K}}+2F_{\max}\sigma^{2}\times\frac{1}{K}\sum_{k=0}^{K-1}\left.\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}\right/\min\{R(W_{k})\}.$
@@ Appendix I Proof of Theorem 3: Convergence of Residual Learning
This section provides the convergence guarantee of the Tiki-Taka under the strongly convex assumption. See 3
@@ I.1 Main proof
The proof of the Tiki-Taka convergence relies on the following two lemmas, which provide the sufficient descent of $W_{k}$ and $\bar{W}_{k}$, respectively.
@@
Suppose Assumptions 12 hold. It holds for Tiki-Taka that
@@
It holds for Tiki-Taka that
The proof of Lemma 5 and 6 are deferred to Section I.2 and I.3, respectively. For a sufficiently large $\gamma$, $P^{*}(W_{k})$ is ensured to be located in the dynamic range of the analog array $P_{k}$. Therefore, we may assume both $q_{+}(P_{k})$ and $q_{-}(P_{k})$ are non-zero, equivalently, there exists a non-zero constant $R^{\texttt{RL}}_{\min}$ such that $\min\{R(P_{k})\}\geq R^{\texttt{RL}}_{\min}$ for all $k$. Under this condition, we have the following inequalities
$\frac{\alpha}{4F_{\max}}\|\nabla f(\bar{W}_{k})\|^{2}_{R(P_{k})}\geq$ $\ \frac{\alpha R^{\texttt{RL}}_{\min}}{4F_{\max}}\|\nabla f(\bar{W}_{k})\|^{2},$ (70)
$\frac{F_{\max}}{\alpha\gamma}\|W_{k+1}-W_{k}\|^{2}_{R(P_{k})^{\dagger}}\leq$ $\ \frac{F_{\max}}{\alpha\gamma R^{\texttt{RL}}_{\min}}\|W_{k+1}-W_{k}\|^{2}.$ (71)
Similarly, we bound the term $\left\|P_{k+1}-P^{*}(W_{k})\right\|^{2}_{R(W_{k})^{\dagger}}$ in (5) by
$\frac{2\beta F_{\max}^{3}}{\gamma}\left\|P_{k+1}-P^{*}(W_{k})\right\|^{2}_{R(W_{k})^{\dagger}}\leq\frac{2\beta F_{\max}^{3}}{\gamma\min\{R(W_{k})\}}\left\|P_{k+1}-P^{*}(W_{k})\right\|^{2}.$ (72)
Notice it is only required to have $\min\{R(W_{k})\}>0$ for the inequality to hold.
@@
$\ \frac{F_{\max}}{\alpha}\mathbb{E}_{\xi_{k}}\left[\|W_{k+1}-W_{k}\|^{2}_{R(P_{k})^{\dagger}}\right]+\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]$ (73)
$=$ $\ \frac{F_{\max}}{\alpha R^{\texttt{RL}}_{\min}}\mathbb{E}_{\xi_{k}}\left[\|W_{k+1}-W_{k}\|^{2}\right]+\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]$
$\overset{(a)}{\leq}$ $\ \frac{2F_{\max}}{\alpha R^{\texttt{RL}}_{\min}}\mathbb{E}_{\xi_{k}}\left[\|W_{k+1}-W_{k}\|^{2}\right]=\frac{2\beta^{2}F_{\max}}{\alpha R^{\texttt{RL}}_{\min}}\|P_{k+1}\odot F(W_{k})-|P_{k+1}|\odot G(W_{k})\|^{2}$
$\leq$ $\ \frac{4\beta^{2}F_{\max}}{\alpha R^{\texttt{RL}}_{\min}}\|P^{*}(W_{k})\odot F(W_{k})-|P^{*}(W_{k})|\odot G(W_{k})\|^{2}$
$\ +\frac{4\beta^{2}F_{\max}}{\alpha R^{\texttt{RL}}_{\min}}\|P_{k+1}\odot F(W_{k})-|P_{k+1}|\odot G(W_{k})-(P^{*}(W_{k})\odot F(W_{k})-|P^{*}(W_{k})|\odot G(W_{k}))\|^{2}$
$\overset{(b)}{\leq}$ $\ \frac{4\beta^{2}F_{\max}}{\alpha R^{\texttt{RL}}_{\min}}\|P^{*}(W_{k})\odot F(W_{k})-|P^{*}(W_{k})|\odot G(W_{k})\|^{2}+\frac{4\beta^{2}F_{\max}}{\alpha R^{\texttt{RL}}_{\min}}\|P_{k+1}-P^{*}(W_{k})\|^{2}.$
where $(a)$ holds if learning rate $\alpha$ is sufficiently small such that $\frac{F_{\max}}{\alpha\gamma R^{\texttt{RL}}_{\min}}\geq 1$; $(b)$ comes from the Lipschitz continuity of the analog update (c.f. Lemma 3).
@@
$\leq$ $\ f(\bar{W}_{k})-f^{*}-\frac{\alpha}{4F_{\max}}\|\nabla f(\bar{W}_{k})\|^{2}_{R(P_{k})}+2\alpha\sigma^{2}\left\|\frac{G(P_{k})}{\sqrt{F(P_{k})}}\right\|^{2}_{\infty}+2\alpha^{2}LF_{\max}^{2}\sigma^{2}$
@@
$\ +\frac{4\beta^{2}F_{\max}}{\alpha R^{\texttt{RL}}_{\min}}\|P^{*}(W_{k})\odot F(W_{k})-|P^{*}(W_{k})|\odot G(W_{k})\|^{2}+\frac{4\beta^{2}F_{\max}}{\alpha R^{\texttt{RL}}_{\min}}{\mathbb{E}}_{\xi_{k}}[\|P_{k+1}-P^{*}(W_{k})\|^{2}]$
$\ +C\Bigg(~\|W_{k}-W^{*}\|^{2}-\frac{\beta}{2\gamma F_{\max}}\|W_{k}-W^{*}\|^{2}_{R(W_{k})}+\frac{3\beta F_{\max}^{3}}{\gamma\min\{R(W_{k})\}}{\mathbb{E}}_{\xi_{k}}[\|P_{k+1}-P^{*}(W_{k})\|^{2}]$
@@
$\leq$ $\ {\mathbb{V}}_{k}-\frac{\alpha R^{\texttt{RL}}_{\min}}{4F_{\max}}\|\nabla f(\bar{W}_{k})\|^{2}+2\alpha\sigma^{2}\left\|\frac{G(P_{k})}{\sqrt{F(P_{k})}}\right\|^{2}_{\infty}+2\alpha^{2}LF_{\max}^{2}\sigma^{2}$
@@
$\ -\left(\frac{\beta\gamma}{2F_{\max}}C-\frac{4\beta^{2}F_{\max}}{\alpha R^{\texttt{RL}}_{\min}}\right)\left\|P^{*}(W_{k})\odot F(W_{k})-|P^{*}(W_{k})|\odot G(W_{k})\right\|^{2}$
$\ +\left(\frac{3\beta F_{\max}^{3}}{\gamma\min\{R(W_{k})\}}C+\frac{4\beta^{2}F_{\max}}{\alpha R^{\texttt{RL}}_{\min}}\right){\mathbb{E}}_{\xi_{k}}[\|P_{k+1}-P^{*}(W_{k})\|^{2}]-\frac{\beta}{2\gamma F_{\max}}C\|W_{k}-W^{*}\|^{2}_{R(W_{k})}.$
Let $C=\frac{10\beta F_{\max}^{2}}{\alpha R^{\texttt{RL}}_{\min}\gamma}$, which leads to the positive coefficient in front of $\|P_{k+1}-P^{*}(W_{k})\|^{2}$, i.e.,
@@
$\leq$ $\ {\mathbb{V}}_{k}-\frac{\alpha R^{\texttt{RL}}_{\min}}{4F_{\max}}\|\nabla f(\bar{W}_{k})\|^{2}+2\alpha\sigma^{2}\left\|\frac{G(P_{k})}{\sqrt{F(P_{k})}}\right\|^{2}_{\infty}+2\alpha^{2}LF_{\max}^{2}\sigma^{2}$
@@
$\ -\frac{\beta^{2}F_{\max}}{\alpha R^{\texttt{RL}}_{\min}}\left\|P^{*}(W_{k})\odot F(W_{k})-|P^{*}(W_{k})|\odot G(W_{k})\right\|^{2}$
$\ +\left(\frac{30\beta^{2}F_{\max}^{5}}{\alpha\gamma\min\{R(W_{k})\}R^{\texttt{RL}}_{\min}}+\frac{4\beta^{2}F_{\max}}{\alpha R^{\texttt{RL}}_{\min}}\right){\mathbb{E}}_{\xi_{k}}[\|P_{k+1}-P^{*}(W_{k})\|^{2}]-\frac{5\beta^{2}F_{\max}}{\alpha R^{\texttt{RL}}_{\min}}\|W_{k}-W^{*}\|^{2}_{R(W_{k})}.$
@@
Suppose Assumptions 1-2 and 4 hold. It holds for Tiki-Taka that

### v5

@@ Appendix H Proof of Theorem 2: Convergence of Analog SGD
Define the saturation vector $M(W_{k})\in{\mathbb{R}}^{D}$ by
$M(W_{k}):=$ $\ {F(W_{k})^{\odot 2}-G(W_{k})^{\odot 2}}={(F(W_{k})+G(W_{k}))\odot(F(W_{k})-G(W_{k}))}$ (65)
@@
$=$ $\ -\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{M(W_{k})}\leq 0.$
@@
$=$ $\ -\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{M(W_{k})}+\frac{\alpha\sigma^{2}}{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$
@@
$\leq$ $\ f(W_{k})-\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{M(W_{k})}+\alpha^{2}LF_{\max}^{2}\sigma^{2}+\frac{\alpha\sigma^{2}}{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$
@@
$\mathbb{E}_{\xi_{k}}[f(W_{k+1})]\leq$ $\ f(W_{k})-\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{M(W_{k})}+\alpha^{2}LF_{\max}^{2}\sigma^{2}+\alpha\sigma^{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$ (72)
@@
$E^{\texttt{ASGD}}_{K}=$ $\ \frac{1}{K}\sum_{k=0}^{K}{\mathbb{E}}[\|\nabla f(W_{k})\odot F(W_{k})+|\nabla f(W_{k})|\odot G(W_{k})\|^{2}+4\|\nabla f(W_{k})\|^{2}_{M(W_{k})}]$ (73)
@@
Assuming the saturation never happens during the training, i.e. $M(W_{k})\geq M^{\texttt{RL}}_{\min}>0$ for all $k\in[K]$, we get a tighter bound in (72) by leveraging $\left\|\nabla f(W_{k})\right\|^{2}_{M(W_{k})}\geq\min\{M(W_{k})\}\left\|\nabla f(W_{k})\right\|^{2}\geq M^{\texttt{RL}}_{\min}\left\|\nabla f(W_{k})\right\|^{2}$
$\mathbb{E}_{\xi_{k}}[f(W_{k+1})]\leq$ $\ f(W_{k})-\frac{\alpha}{2F_{\max}}\|\nabla f(W_{k})\|^{2}_{M(W_{k})}+\alpha^{2}LF_{\max}^{2}\sigma^{2}+\alpha\sigma^{2}\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}$ (74)
@@
$=$ $\ \frac{4F_{\max}^{2}}{M^{\texttt{RL}}_{\min}}\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L}{K}}+2F_{\max}\sigma^{2}\times\frac{1}{K}\sum_{k=0}^{K-1}\left.\left\|\frac{G(W_{k})}{\sqrt{F(W_{k})}}\right\|^{2}_{\infty}\right/\min\{M(W_{k})\}.$
@@ Appendix I Proof of Theorem 3: Convergence of Residual Learning
This section provides the convergence guarantee of the Residual Learning under the strongly convex assumption. See 3
@@ I.1 Main proof
The proof relies on the following two lemmas, which provide the sufficient descent of $W_{k}$ and $\bar{W}_{k}$, respectively.
@@
Suppose Assumptions 12 hold. It holds for Residual Learning that
@@
It holds for Residual Learning that
The proof of Lemma 5 and 6 are deferred to Section I.2 and I.3, respectively. For a sufficiently large $\gamma$, $P^{*}(W_{k})$ is ensured to be located in the dynamic range of the analog array $P_{k}$. Therefore, we may assume both $q_{+}(P_{k})$ and $q_{-}(P_{k})$ are non-zero, equivalently, there exists a non-zero constant $M^{\texttt{RL}}_{\min}$ such that $\min\{M(P_{k})\}\geq M^{\texttt{RL}}_{\min}$ for all $k$. Under this condition, we have the following inequalities
$\frac{\alpha}{4F_{\max}}\|\nabla f(\bar{W}_{k})\|^{2}_{M(P_{k})}\geq$ $\ \frac{\alpha M^{\texttt{RL}}_{\min}}{4F_{\max}}\|\nabla f(\bar{W}_{k})\|^{2},$ (76)
$\frac{F_{\max}}{\alpha\gamma}\|W_{k+1}-W_{k}\|^{2}_{M(P_{k})^{\dagger}}\leq$ $\ \frac{F_{\max}}{\alpha\gamma M^{\texttt{RL}}_{\min}}\|W_{k+1}-W_{k}\|^{2}.$ (77)
Similarly, we bound the term $\left\|P_{k+1}-P^{*}(W_{k})\right\|^{2}_{M(W_{k})^{\dagger}}$ in (5) by
$\frac{2\beta F_{\max}^{3}}{\gamma}\left\|P_{k+1}-P^{*}(W_{k})\right\|^{2}_{M(W_{k})^{\dagger}}\leq\frac{2\beta F_{\max}^{3}}{\gamma\min\{M(W_{k})\}}\left\|P_{k+1}-P^{*}(W_{k})\right\|^{2}.$ (78)
Notice it is only required to have $\min\{M(W_{k})\}>0$ for the inequality to hold.
@@
$\ \frac{F_{\max}}{\alpha}\mathbb{E}_{\xi_{k}}\left[\|W_{k+1}-W_{k}\|^{2}_{M(P_{k})^{\dagger}}\right]+\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]$ (79)
$=$ $\ \frac{F_{\max}}{\alpha M^{\texttt{RL}}_{\min}}\mathbb{E}_{\xi_{k}}\left[\|W_{k+1}-W_{k}\|^{2}\right]+\mathbb{E}_{\xi_{k}}[\|W_{k+1}-W_{k}\|^{2}]$
$\overset{(a)}{\leq}$ $\ \frac{2F_{\max}}{\alpha M^{\texttt{RL}}_{\min}}\mathbb{E}_{\xi_{k}}\left[\|W_{k+1}-W_{k}\|^{2}\right]=\frac{2\beta^{2}F_{\max}}{\alpha M^{\texttt{RL}}_{\min}}\|P_{k+1}\odot F(W_{k})-|P_{k+1}|\odot G(W_{k})\|^{2}$
$\leq$ $\ \frac{4\beta^{2}F_{\max}}{\alpha M^{\texttt{RL}}_{\min}}\|P^{*}(W_{k})\odot F(W_{k})-|P^{*}(W_{k})|\odot G(W_{k})\|^{2}$
$\ +\frac{4\beta^{2}F_{\max}}{\alpha M^{\texttt{RL}}_{\min}}\|P_{k+1}\odot F(W_{k})-|P_{k+1}|\odot G(W_{k})-(P^{*}(W_{k})\odot F(W_{k})-|P^{*}(W_{k})|\odot G(W_{k}))\|^{2}$
$\overset{(b)}{\leq}$ $\ \frac{4\beta^{2}F_{\max}}{\alpha M^{\texttt{RL}}_{\min}}\|P^{*}(W_{k})\odot F(W_{k})-|P^{*}(W_{k})|\odot G(W_{k})\|^{2}+\frac{4\beta^{2}F_{\max}}{\alpha M^{\texttt{RL}}_{\min}}\|P_{k+1}-P^{*}(W_{k})\|^{2}.$
where $(a)$ holds if learning rate $\alpha$ is sufficiently small such that $\frac{F_{\max}}{\alpha\gamma M^{\texttt{RL}}_{\min}}\geq 1$; $(b)$ comes from the Lipschitz continuity of the analog update (c.f. Lemma 3).
@@
$\leq$ $\ f(\bar{W}_{k})-f^{*}-\frac{\alpha}{4F_{\max}}\|\nabla f(\bar{W}_{k})\|^{2}_{M(P_{k})}+2\alpha\sigma^{2}\left\|\frac{G(P_{k})}{\sqrt{F(P_{k})}}\right\|^{2}_{\infty}+2\alpha^{2}LF_{\max}^{2}\sigma^{2}$
@@
$\ +\frac{4\beta^{2}F_{\max}}{\alpha M^{\texttt{RL}}_{\min}}\|P^{*}(W_{k})\odot F(W_{k})-|P^{*}(W_{k})|\odot G(W_{k})\|^{2}+\frac{4\beta^{2}F_{\max}}{\alpha M^{\texttt{RL}}_{\min}}{\mathbb{E}}_{\xi_{k}}[\|P_{k+1}-P^{*}(W_{k})\|^{2}]$
$\ +C\Bigg(~\|W_{k}-W^{*}\|^{2}-\frac{\beta}{2\gamma F_{\max}}\|W_{k}-W^{*}\|^{2}_{M(W_{k})}+\frac{3\beta F_{\max}^{3}}{\gamma\min\{M(W_{k})\}}{\mathbb{E}}_{\xi_{k}}[\|P_{k+1}-P^{*}(W_{k})\|^{2}]$
@@
$\leq$ $\ {\mathbb{V}}_{k}-\frac{\alpha M^{\texttt{RL}}_{\min}}{4F_{\max}}\|\nabla f(\bar{W}_{k})\|^{2}+2\alpha\sigma^{2}\left\|\frac{G(P_{k})}{\sqrt{F(P_{k})}}\right\|^{2}_{\infty}+2\alpha^{2}LF_{\max}^{2}\sigma^{2}$
@@
$\ -\left(\frac{\beta\gamma}{2F_{\max}}C-\frac{4\beta^{2}F_{\max}}{\alpha M^{\texttt{RL}}_{\min}}\right)\left\|P^{*}(W_{k})\odot F(W_{k})-|P^{*}(W_{k})|\odot G(W_{k})\right\|^{2}$
$\ +\left(\frac{3\beta F_{\max}^{3}}{\gamma\min\{M(W_{k})\}}C+\frac{4\beta^{2}F_{\max}}{\alpha M^{\texttt{RL}}_{\min}}\right){\mathbb{E}}_{\xi_{k}}[\|P_{k+1}-P^{*}(W_{k})\|^{2}]-\frac{\beta}{2\gamma F_{\max}}C\|W_{k}-W^{*}\|^{2}_{M(W_{k})}.$
Let $C=\frac{10\beta F_{\max}^{2}}{\alpha M^{\texttt{RL}}_{\min}\gamma}$, which leads to the positive coefficient in front of $\|P_{k+1}-P^{*}(W_{k})\|^{2}$, i.e.,
@@
$\leq$ $\ {\mathbb{V}}_{k}-\frac{\alpha M^{\texttt{RL}}_{\min}}{4F_{\max}}\|\nabla f(\bar{W}_{k})\|^{2}+2\alpha\sigma^{2}\left\|\frac{G(P_{k})}{\sqrt{F(P_{k})}}\right\|^{2}_{\infty}+2\alpha^{2}LF_{\max}^{2}\sigma^{2}$
@@
$\ -\frac{\beta^{2}F_{\max}}{\alpha M^{\texttt{RL}}_{\min}}\left\|P^{*}(W_{k})\odot F(W_{k})-|P^{*}(W_{k})|\odot G(W_{k})\right\|^{2}$
$\ +\left(\frac{30\beta^{2}F_{\max}^{5}}{\alpha\gamma\min\{M(W_{k})\}M^{\texttt{RL}}_{\min}}+\frac{4\beta^{2}F_{\max}}{\alpha M^{\texttt{RL}}_{\min}}\right){\mathbb{E}}_{\xi_{k}}[\|P_{k+1}-P^{*}(W_{k})\|^{2}]-\frac{5\beta^{2}F_{\max}}{\alpha M^{\texttt{RL}}_{\min}}\|W_{k}-W^{*}\|^{2}_{M(W_{k})}.$
@@
Suppose Assumptions 1-2 and 3 hold. It holds for Tiki-Taka that